A certain charge Q is divided at first into 2 parts q &( Q-q) later on the charges are placed at a certain distance. If the force of interaction between the two charges is maximum then Q/q is

Grade:11

1 Answers

Apptica Corporations
10 years ago
Answer-2 Since here charge `q` can be taken as a variable Force= K q*(Q-q)/(r^2) For Force to be maximum, q*(Q-q) is maximum Differentiating w.r.t. q df/dq= Q-2q df/dq=0 Q/q=2 To check more you can calculate double differentiation of force f. If its negative then for that value of q you get maximum force